Show your work:
How many gallons of rocket fuel containing 80% liquid oxygen must be mixed with 250 gallons containing 40% liquid oxygen to produce a mixture containing 70% liquid oxeygen

Answer by checkley79(3341)

.80X+.40*250=.70(250+X)
.80X+100=175+.70X
.80X-.70X=175-100
.10X=75
X=75/.10
X=750 GAL OF 80% OXYGEN IS USED.
PROOF:
.80*750+.40*250=.70(250+750)
600+100=.7*1000
700=700